OpenStudy (anonymous):

The base of a pyramid is a rectangle with a width of 6 cm and a length of 8 cm . Find , in centimeters , the height of the pyramid if the volume is 288 cm^3.

OpenStudy (anonymous):

volume = 1/3 x b x h 288 = 1/3 x 6(8) x h 288 = 16h 288/16 = h 18cm = h

OpenStudy (anonymous):

thank you seashell!

OpenStudy (anonymous):

no prob :)